Based on economists’ forecasts and analysis, 1-year Treasury bill rates and liquidity premiums for the next four years are expected to be as follows:

R1 = 1.85 %
E(2r1) = 2.75 % L2 = 0.06 %
E(3r1) = 3.15 % L3 = 0.08 %
E(4r1) = 3.60 % L4 = 0.13 %

Using the liquidity premium theory, determine the current (long-term) rates. (Do not round intermediate calculations. Round your answers to 2 decimal places.)

Years/Current (long-Term) Rates

1

2

3

4
Answer #1

The current (long term) rates are calculated as below:

Year 1 = R1 = 1.85%

Year 2 = [(1+R1 )*(1+E(2r1)+L2)]*(1/2) - 1 = [(1+1.85%)*(1+2.75%+0.06%)]*(1/2) - 1 = 2.33%

Year 3 = [(1+R1 )*(1+E(2r1)+L2)*(1+E(3r1)+L3)]*(1/3) - 1 = [(1+1.85%)*(1+2.75%+0.06%)*(1+3.15%+0.08%)]^(1/3) - 1 = 2.63%

Year 4 = [(1+R1 )*(1+E(2r1)+L2)*(1+E(3r1)+L3)*(1+E(4r1)+L4)]*(1/4) - 1 = [(1+1.85%)*(1+2.75%+0.06%)*(1+3.15%+0.08%)*(1+3.60%+0.13%)]^(1/4) - 1 = 2.90%

_____

Tabular Representation:

Years Current (long term) Rates
1 1.85%
2 2.33%
3 2.63%
4 2.90%